OpenStudy (anonymous):

Provide a complete solution with all work shown. Each person in a random sample of 1,026 adults in the United States was asked the following question: "Based on what you know about the Social Security system today, what would you like Congress and the President to do during this next year?" The response choices and the percentages selecting them are shown below: Completely overhaul the system 19% Make some major changes 39% Make some minor adjustments 30% Leave the system the way it is now 11% No opinion 1%

OpenStudy (anonymous):

a. Find a 95% confidence interval for the proportion of all United States adults who would respond "Make some major changes" to the question. Give an interpretation of the confidence interval and give an interpretation of the confidence level. b. An advocate for leaving the system as it is now commented. "Based on this poll, only 39% of adults in the sample responded that they want some major changes made to the system, while 41% responded that they want only minor changes or no changes at all. Therefore, we should not change the system." Explain why this statement, while technically correct, is misleading.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@h

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

the idea of this site is to NOT show all the work ... but to get you to participate in the solution process.

OpenStudy (amistre64):

a confidence interval will assume that the stated percentage is accurate; and give a range wherein we believe the actual population statisitic rests.

OpenStudy (amistre64):

as such, we need to address:\[\hat p \pm Z_{.99}(\sigma/\sqrt n)\]

OpenStudy (anonymous):

I understand that is not the purpose of the site, I just copied and pasted the entire question, and that was part of it.[But if you wanted to give me the answer I wouldn't object. ;-)]

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

i would object ;)

OpenStudy (amistre64):

do you know the zscore related to 99% about the mean? might be read in someplaces as an alpha value of a 2-tailed thing ... 1/2% = .005

OpenStudy (amistre64):

assuming memory holds: z = 2.576 was in my book way back when the portion they are seeking for look st o be associated with .39 in favor using this as the assumed correct parametric, we can assess the standard deviation as sqrt(.39(1-.39)) ... sqrt(pq) in some texts ... adjusted by a division of sqrt(n), n being the sample size

OpenStudy (amistre64):

as such, we can be 99% confident that the actual population statistic is somplace between:\[.39-2.576\sqrt{\frac{.39(.61)}{1026}}~and~~.39+2.576\sqrt{\frac{.39(.61)}{1026}}\]

OpenStudy (amistre64):

i got no good ideas for the part b tho ... maybe the guy is a democratic?

OpenStudy (amistre64):

*democrat

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

The person who was previously helping me is no longer online.

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Is the interval (.3480,.8230)?